40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 3 01:49:46 2013, Annalisa, Update, Green Locking, ALS servo configuration ALSALS_calibrated
    Reply  Wed Jul 3 03:06:29 2013, Annalisa, Update, Green Locking, ALS servo configuration 
Message ID: 8792     Entry time: Wed Jul 3 01:49:46 2013     Reply to this: 8793  
Author: Annalisa 
Type: Update 
Category: Green Locking 
Subject: ALS servo configuration 

[Koji, Annalisa, Manasa]

Today we worked on the ALS servo stabilization for the Y arm.

First step: find the beat note

The beat note was found following the usual steps:

  • Y arm cavity locked on IR to have a good alignment
  • Y arm cavity locked on green (eventually unlocked on IR)
  • beat note alignment maximized on the PSL table

Beat note amplitude = -27 dBm @ 50 MHz

PSL temperature = 31.54 degC

Laser Offset on the slow servo2 = -11011

 

In the GREEN HORNET we did the following changes for the Y arm:

Input Signal Conditioning

On the C1ALS-BEATY_FINE  screen the same antiwhitening filters of the C1ALS-BEATX_FINE have been reproduced. At moment, only the FM3 [10:1] is enabled.

On the C1ALS-BEATY_FINE_PHASE screen the gain was set at 3600, since the amplitude of the Q signal after the Phase rotator (BEATY_FINE_Q_ERR) was about 30. To set this value we made a proportion with respect to a previous optimized value, where the amplitude was 100 and the gain was set to 1200.

DOF filters

In order to stabilize the beat frequency, we started enabling the FM5 [1000:1] filter in the C1ALS_YARM panel, and then we started increasing the gain first in small steps (0.1), in order to understand which sign the gain should have without kicking the mirror.

We measured the Power Spectrum of the C1:ALS-BEATY_FINE_PHASE_OUT in-loop signal while varying the gain of the C1ALS_YARM servo filter.

Eventually, we enabled the following filters:

FM2 [0:1]

FM3 [1:5]

FM4 [1:50]

FM5 [1000:1]

FM6 [RG3.2]

FM7 [RG16.5]

Gain = -30.

Koji expects the UGF of the loop to be around 100-ish Hz, and he also expected the small bump around 300-400 Hz.

Then we realized that the channel we were measuring was not calibrated in unit of Hz, so we took again the measurement looking at the channel C1:ALS-BEATY_FINE_PHASE_OUT_HZ. In this case, we didn't observe any bump. Maybe the beat frequency was slightly changed from the previous measurement and the all servo shape was also different. The final value of the gain was set at -8.

The Y axis unit is missing (bad me!). It's in deg/sqrt(Hz) for the first plot and Hz/sqrt(Hz) for the second one.

 

Attachment 1: ALS  22 kB  Uploaded Wed Jul 3 15:17:37 2013
Attachment 2: ALS_calibrated  24 kB  Uploaded Wed Jul 3 15:17:47 2013
ELOG V3.1.3-